Circular No. 7068

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

CATACLYSMIC VARIABLE IN AURIGA J. Y. Wei, Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astronomical

Observatory (BAO), report their discovery in the course of the BAO supernova survey of an outburst of the cataclysmic variable noted on IAUC 6758. The unfiltered CCD magnitudes of the star were 16.6 on Dec. 9.73, 16.0 on Dec. 12.73, and 16.6 on Dec. 14.78 UT. The star was not present on a CCD image obtained on Dec. 3.

SUPERNOVA 1997eg IN NGC 5012 C. K. Lacey and K. W. Weiler, Naval Research Laboratory; R. A.

Sramek, National Radio Astronomy Observatory (NRAO); and S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology, report the detection of radio emission from the type-IIn supernova 1997eg (IAUC 6790, 6794, 6823, 6917) with the Very Large Array: “A radio-flux density of 0.52 ± 0.06 mJy at 8.44 GHz (wavelength 3.6 cm) on May 31 and 0.53 ± 0.12 mJy at 8.46 GHz on June 9 has been measured at a position of α = 13^{h}11^{m}36^{s}.71, δ = +22^{o}55′28′′.9 (equinox 2000.0; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (end figures 36^{s}.73, 29′′.4; IAUC 6790). Nondetections of < 0.21 mJy (3σ) at 8.44 GHz (wavelength 3.6 cm) and < 0.84 mJy (3σ) at 14.96 GHz (wavelength 2.0 cm) on Feb. 10, < 0.17 mJy (3σ) at 4.89 GHz (wavelength 6.0 cm) on Feb. 13, < 0.44 mJy (3σ) at 14.96 GHz (wavelength 2.0 cm) on May 31, and < 0.60 mJy (3σ) at 14.96 GHz (wavelength 2.0 cm) on June 9 have also been obtained. Filippenko and Barth (IAUC 6794) predict significant radio and x-ray emission from SN 1997eg based on comparison of its optical spectral characteristics with the class of luminous radio and x-ray emitting supernovae resembling SN 1988Z. If the Filippenko-Barth prediction is correct, SN 1997eg could reach a maximum 6-cm flux density of > 10 mJy, assuming a distance of 40 Mpc (Ho et al. 1997, Ap.J. Suppl. 12, 315), as late as four years after explosion. The above detections in May–June imply that SN 1997eg was significantly more radio luminous at 8.46 GHz by age ∼ 180 days than the best-fit model for the radio emission from SN 1988Z would predict. We encourage further multi-wavelength observations.”

1998 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green